OpenStudy (anonymous):

What is the Greatest Common Factor of x^6 and x^4?

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

whoa dont think i know this one

OpenStudy (anonymous):

i dont either :d

OpenStudy (karatechopper):

ok start by factoring x^6 out

OpenStudy (anonymous):

x^4

OpenStudy (karatechopper):

:/

OpenStudy (karatechopper):

guess i can't help on this anymore

OpenStudy (anonymous):

how did u find it?

OpenStudy (anonymous):

oh now its coming back to me

OpenStudy (anonymous):

x^6=x*x*x*x*x*x x^4=x*x*x*x So x^4 is common